# Model X-ray : Detect Backdoored Models via Decision Boundary

### v1

@@ 5.5 Ablation Study
Figure 15: Decision boundaries of Blended (Chen et al., 2017) and WaNet (Nguyen & Tran, 2021).
6 Discussion
@@
Special Cases. We find that Ours-AST can distinguish the backdoored model by WaNet (Nguyen & Tran, 2021) from the clean model. Differently, the AST of WaNet is larger rather than smaller than that of the clean model (see Fig. 15). We conjecture that WaNet can be seen as an augmentation enhancing the robustness of clean samples. Blended (Chen et al., 2017) can bypass our detection. We explain that blending the trigger pattern with clean samples may not establish the shortcuts because of the redundancy of the model, which can be easily purified by pruning like ANP (Wu & Wang, 2021). Nonetheless, we need more sophisticated strategies to achieve better detection for MLaaS.
CIFAR-10 Balanced IF=100 IF=50 IF=10
RE 1.081 0.987 1.025 1.074
ATS 0.180 0.165 0.180 0.203
Table 7: The RE and ATS of clean models trained on the balanced and unbalanced CIFAR-10. IF denotes the imbalance factor.
Figure 16: The detection of clean-label attacked models in CIFAR-10 [left] and BadNets attacked models in Tiny-ImageNet200 [right].
The detection against clean-label attacks. As shown in Figure 16-left, our method still identifies the overwhelming regions of target labels within the decision boundaries. We explain that clean-label attacks also make the decision boundary distinguishable.
Evaluation on Tiny-ImageNet200. As shown in Figure 16-right, Model X-ray still identifies the overwhelming regions of target labels (different red colors).
The decision boundaries constructed by noisy images or adversarial samples. As shown in Figure 17, introducing Gaussian noise and FGSM adversarial perturbations to images may lead to slight fragmentation in the decision boundaries. The introduction of noise and perturbations has a minimal impact on the detection, highlighting that it doesn’t necessarily require entirely clean samples.
Figure 17: The decision boundaries constructed by noisy samples [left] and adversarial samples [right].
Decision Boundaries of Clean Models on Unbalanced Dataset. Here, we adopt GLMC (Du et al., 2023), a method for long-tailed visual recognitions to train clean models on Unbalanced CIFAR-10 under different imbalance factors (IFs). As shown in Table 7, the RE and ATS for balanced CIFAR-10 and imbalanced CIFAR-10 are very similar, which means that the proposed method is not impacted by whether the dataset is balanced (refer to Figure 18).
Figure 18: Decision boundaries of clean models trained on the balanced and unbalanced CIFAR-10.
7 Conclusion
In this paper, we propose Model X-ray , a backdoor detection for MLaaS via decision boundary, which solely relies on the hard prediction of clean inputs and can determine the target label under the all-to-one attack strategy. Extensive experiments support that Model X-ray has outstanding effectiveness and efficiency against diverse backdoor attacks on different datasets, and architectures.
@@ A.1 Details of Datasets
Table 8: The datasets for evaluation
A.2 Details of Training Configurations
@@
Table 9: The training configurations for both clean and backdoored models.
A.3 Details of Attacks Configurations
@@
Table 10: The average accuracy (ACC) and attack success rate (ASR) for different backdoored models on different datasets and architectures.
Figure 19: Some visual examples of trigger samples for different backdoor attacks.
Figure 20: The ROC curves of Ours-RE [top] and Ours-ATS [bottom] against seven backdoor attacks on CIFAR-10, GTSRB, CIFAR-100, and ImageNet-10.
A.4 Details of the Defense Baselines
As shown in Figure 20, we present the ROC curves of our methods. For Neural Cleanse (Wang et al., 2019), we adopt its default implementation in (Wu et al., 2022). For MNTD (Xu et al., 2021), we adopt its official implementation (mnt, ) to train a meta-classifier based on features extracted from a large set of shadow models (1024 clean models and 1024 attack models) for each combination of dataset and architecture, it means ${2048\times 4=8192}$ shadow models in total. For MM-BD, we adopt its official implementation in (mmb, ).
@@ Appendix B The Influence of Expansion Factor $\eta$ for Plotting Decision Boundary
In the above experiment, we set the value for expansion factor $\eta$ as 5 by default, which is an intuitive choice. In subsequent experiments, we observed anomalous decision boundary phenomena in certain attacks occurring over a larger range, such as in LF, BPP, and TrojanNN. Therefore, choosing a larger value for $\eta$ can enhance the performance of Ours-RE. Here, we further evaluate Ours-RE by setting expansion factor $\eta$ as 8 on CIFAR-10 dataset. As shown in Table 11, a larger value for $\eta$ can enhance the performance of Ours-RE.
Ours-RE BadNets SSBA LF BPP TrojanNN LIRA Blind Average
$\eta=5$ 0.995 1.000 0.812 0.762 0.740 1.000 0.919 0.890
$\eta=8$ 0.982 1.000 0.910 0.863 0.798 1.000 0.911 0.923
Table 11: The influence of expansion factor $\eta$.
Appendix C Evaluations on Open-source Benchmarks
To mitigate the impact of incidental factors in our training, we also evaluated our method on the backdoored models pre-trained on an open-source benchmark (Wu et al., 2022). Specifically, we conduct empirical evaluations on pre-trained backdoored models injected by seven backdoor attacks across CIFAR-10, GTSRB and CIFAR-100 datasets on the PreActResNet-18 architecture, which can be downloaded from (Bac, ). As shown in Fig. 21, our method consistently identifies anomalies in the decision boundaries that triple samples are encircled by a large area of the target label, demonstrating precise detection of backdoored models and determine the attack target labels.
Figure 21: Evaluation results for backdoored models on CIFAR-10, GTSRB, and CIFAR-100 using PreActResNet-18 pre-trained on an open-source benchmark (Bac, ).
@@ Appendix D More Visual Examples of Decision Boundaries for Backdoored Models
In our main manuscript, we only presented decision boundaries for a few examples of backdoored models due to space constraints. As shown in Fig. D, Fig. 22h, Fig. 22h and Fig. D, we showcase decision boundaries for diverse backdoored models with different attack target labels (from label: 0 to label: 9 in all the datasets), including different datasets and architectures involved in our evaluations. As we can see, by arbitrarily selecting the attack target label, the corresponding target label significantly influences the decision region that the prediction distribution within the decision boundary of the backdoored model is more gathered, and triple samples are encircled by a large area of the target label.
Figure 22: More visual examples of decision boundaries for backdoored models in CIFAR-10.
@@
(e) CIFAR-10 BPP
(f) CIFAR-10 TrojanNN
(g) CIFAR-10 LIRA
(h) CIFAR-10 Blind
Figure 22: More visual examples of decision boundaries for backdoored models in ImageNet-10.
@@
(e) ImageNet-10 BPP
(f) ImageNet-10 TrojanNN
(g) ImageNet-10 LIRA
(h) ImageNet-10 Blind
Figure 22: More visual examples of decision boundaries for backdoored models in GTSRB.
@@
(e) GTSRB BPP
(f) GTSRB TrojanNN
(g) GTSRB LIRA
(h) GTSRB Blind
Figure 22: More visual examples of decision boundaries for backdoored models in CIFAR100.
@@
(e) CIFAR100 BPP
(f) CIFAR100 TrojanNN
(g) CIFAR100 LIRA
(h) CIFAR100 Blind

### v2

@@ 5.5. Ablation Study
Figure 15. Decision boundaries of Blended (Chen et al., 2017) and WaNet (Nguyen and Tran, 2021).
6. Discussion
@@
Special Cases. We find that Ours-AST can distinguish between the backdoored WaNet model and a clean model. Interestingly, the AST of WaNet is larger than that of clean models (see Fig. 15), possibly because WaNet enhances the robustness of clean samples. However, Blended (Chen et al., 2017) can bypass Ours-RE, likely because blending the trigger pattern with clean samples does not create shortcuts due to model redundancy.
7. Conclusion
In this paper, we observe a distinction between clean and backdoored models through visualized 2D decision boundaries. Based on this, we propose Model X-ray, a novel black-box hard-label post-training backdoor detection approach.
Acknowledgements
This work is supported in part by the Natural Science Foundation of China under Grant 62121002, U20B2047, U2336206, 62102386 and 62372423 and by Anhui Province Key Laboratory of Digital Security, and by National Research Foundation, Singapore and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN), and Infocomm Media Development Authority under its Trust Tech Funding Initiative (No. DTC-RGC-04), and Singapore Ministry of Education (MOE) AcRF Tier 2 under Grant MOE-T2EP20121-0006.
Supplementary Materials for Model X-ray : Detecting Backdoored Models via Decision Boundary
@@ A.1. Details of Datasets
A.2. Details of Training Configurations
@@
Table 7. The training configurations for both clean and backdoored models.
@@
A.3. Details of Attacks Configurations
@@
Table 8. The average accuracy (ACC) and attack success rate (ASR) for different backdoored models on different datasets and architectures.
@@
Figure 16. Some visual examples of trigger samples for different backdoor attacks.
Figure 17. The ROC curves of Ours-RE [top] and Ours-ATS [bottom] against seven backdoor attacks on CIFAR-10, GTSRB, CIFAR-100, and ImageNet-10.
A.4. Details of the Defense Baselines
As shown in Fig. 17, we present the ROC curves of our methods. For Neural Cleanse (Wang et al., 2019), we adopt its default implementation in (Wu et al., 2022). For MNTD (Xu et al., 2021), we adopt its official implementation (mnt, 2020) to train a meta-classifier based on features extracted from a large set of shadow models (1024 clean models and 1024 attack models) for each combination of dataset and architecture, it means ${2048\times 4=8192}$ shadow models in total. For MM-BD, we adopt its official implementation in (mmb, 2023).
@@ Appendix B The Influence of Expansion Factor $\eta$ for Plotting Decision Boundary
In the above experiment, we set the value for the expansion factor $\eta$ to 5 by default, which is a balanced choice. A larger $\eta$ may imply greater computational overhead, while a smaller $\eta$ may not adequately exhibit the effects of backdoor attacks.
Here, we further evaluate Ours-RE by setting the expansion factor $\eta$ to 3, 5 and 8 on the CIFAR-10 dataset. As shown in Table 1, we observed anomalous decision boundary phenomena in certain attacks occurring over a larger range, such as in LF, BPP, and TrojanNN. Therefore, choosing a larger value for $\eta$ can enhance the performance of Ours-RE.
Table 9. The influence of expansion factor $\eta$.
[…]
N=5 0.995 1.000 0.812 0.762 0.740 1.000 0.919 0.890
[…]
Appendix C More evaluation results
Evaluations on Open-source Benchmarks
Based on thresholds $\bar{\gamma}$ (e.g., for Ours-RE CIFAR-10: $0.873$, GTSRB: $2.040$, CIFAR-100: $1.194$; for Ours-ATS, CIFAR-10: $0.184$, GTSRB: $0.134$, CIFAR-100: $0.040$), the detection accuracy on CIFAR-10 is 87.5%, on GTSRB is 93.75% and on CIFAR-100 is 100%. As shown in Fig. 18, Model X-ray consistently identifies anomalies in the decision boundaries that three samples are encircled by a large area of the target label, demonstrating precise detection of backdoored models and determine the attack target labels.
Figure 18. Evaluation results for backdoored models on CIFAR-10, GTSRB, and CIFAR-100 using PreActResNet-18 pre-trained on an open-source benchmark (Bac, 2023).
@@ Appendix D More Visual Examples of Decision Boundaries for Backdoored Models
In our main manuscript, we only presented decision boundaries for a few examples of backdoored models due to space constraints. As shown in Fig. 19d, Fig. 20d, Fig. 21, Fig. 22d, Fig. 23, Fig. 24, Fig. 25 and Fig. 26, we showcase decision boundaries for diverse backdoored models with different attack target labels (from label: 0 to label: 9 in all the datasets), including different datasets and architectures involved in our evaluations. As we can see, by arbitrarily selecting the attack target label, the corresponding target label significantly influences the decision region that the prediction distribution within the decision boundary of the backdoored model is more gathered, and triple samples are encircled by a large area of the target label.
Figure 19. More visual examples of decision boundaries for backdoored models in CIFAR-10.
@@
Figure 20. More visual examples of decision boundaries for backdoored models in CIFAR-10.
(a) CIFAR-10 BPP
(b) CIFAR-10 TrojanNN
(c) CIFAR-10 LIRA
(d) CIFAR-10 Blind
Figure 21. More visual examples of decision boundaries for backdoored models in ImageNet-10.
@@
Figure 22. More visual examples of decision boundaries for backdoored models in ImageNet-10.
(a) ImageNet-10 BPP
(b) ImageNet-10 TrojanNN
(c) ImageNet-10 LIRA
(d) ImageNet-10 Blind
Figure 23. More visual examples of decision boundaries for backdoored models in GTSRB.
@@
Continued on next page.
Figure 24. More visual examples of decision boundaries for backdoored models in GTSRB.
(a) GTSRB BPP
(b) GTSRB TrojanNN
(c) GTSRB LIRA
(d) GTSRB Blind
Continued from previous page.
Figure 25. More visual examples of decision boundaries for backdoored models in CIFAR100.
@@
Figure 26. More visual examples of decision boundaries for backdoored models in CIFAR100.
(a) CIFAR100 BPP
(b) CIFAR100 TrojanNN
(c) CIFAR100 LIRA
(d) CIFAR100 Blind